OpenStudy (anonymous):

I WILL MEDAL!!!!!!!!!! -7/12 - -1/6

RhondaSommer (rhondasommer):

I can help :)

OpenStudy (anonymous):

okay

rebeccaxhawaii (rebeccaxhawaii):

- and - makes a +

RhondaSommer (rhondasommer):

okay, so first you need to make the denominator the same number so lets do this \[-\frac{ 1 }{6}*\frac{2}{2}\]

RhondaSommer (rhondasommer):

I got this rebecca

OpenStudy (anonymous):

okay

rebeccaxhawaii (rebeccaxhawaii):

so -7/12+1/6

rebeccaxhawaii (rebeccaxhawaii):

and okay @RhondaSommer ill let you have this one

OpenStudy (anonymous):

-7/12 and -2/12

OpenStudy (anonymous):

is it -5/12?

RhondaSommer (rhondasommer):

perfect. now when you subtract a negative you are actually adding a positive. so think of "Stay Flip Flip" so -7/12 (stay) minus (flip so now it is add) -2/12 (flip so now it is just 2/12) \[-\frac{7}{12}+\frac{2}{12}\] which would be \[-\frac{5}{12}\]

RhondaSommer (rhondasommer):

so yes you are correct.

OpenStudy (anonymous):

Thank you so much!